I'm a little fuzzy on this question and could use some expertise:
Let f(x) = 7x^2 + 5 and g(x) = x – 3.
(a) Find the composite function (f o g)(x) and simplify.
(b) Find (fog)(2)
am I writing (7x^2 +5)(x-3)(x) for A and (7x^2 +5)(x-3)(2) for B?
I really appreciate any assistance! Thank you!

Answer by stanbon(75887)

Let f(x) = 7x^2 + 5 and g(x) = x – 3.
(a) Find the composite function (f o g)(x) and simplify.
----
fog(x) = f[g(x)]
= f[x-3] = 7(x-3)^2+5 = 7(x^2-6x+9)+5 = 7x^2 - 42x +68
------------------------------
(b) Find (fog)(2)
fog(2) = f[g(2)] = f[-1] = 7(-1)^2+5 = 12
--------------------------------------------
Cheers,
Stan H.